OpenStudy (anonymous):

(3/5)2 ÷25 please help

OpenStudy (amistre64):

6/125 if i see it good

OpenStudy (amistre64):

9/625 is what myin was wanting to type

OpenStudy (anonymous):

9/625

OpenStudy (anonymous):

\[\frac{3}{5}*\frac{3}{5}*\frac{25} = \frac{9}{625}\]

OpenStudy (anonymous):

does't any one ever work?

myininaya (myininaya):

\[(\frac{3}{5})^2 \cdot \frac{1}{25}=\frac{3^2}{5^2 \cdot 5^2}=\frac{9}{5^4}=\frac{9}{625}\]

OpenStudy (amistre64):

thanx obama!!

OpenStudy (anonymous):

(6/25) * (1/25) =6/625

OpenStudy (anonymous):

Lol got latex wrong

OpenStudy (anonymous):

lol

myininaya (myininaya):

i got 30 minutes of sleep today study for a presentation all night and morning i'm sleepy

OpenStudy (amistre64):

.... is that sulphur i smell?

OpenStudy (anonymous):

have a beer and a nap

myininaya (myininaya):

yes

OpenStudy (amistre64):

lol ;)

OpenStudy (anonymous):

what did present?

myininaya (myininaya):

i have one more class to teach then i will drive the hour home and go to sleep

OpenStudy (anonymous):

not more integrals i hope

myininaya (myininaya):

trig

OpenStudy (amistre64):

i hopes its in irish trig

OpenStudy (anonymous):

trig! great

OpenStudy (anonymous):

aka trickonometry

myininaya (myininaya):

i love right triangles they complete me

OpenStudy (anonymous):

really? squares complete me. or is it the other way around?

myininaya (myininaya):

it depends i'm half a square so a triangle completes me

OpenStudy (anonymous):

enjoy your class. i am done. almost beer o'clock